OpenStudy (anonymous):

6 is greater than or equal to -4 over 5n

zepdrix (zepdrix):

Hi there blake! \(\Large \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\) \[\Large 6\quad\gt\quad -\frac{4}{5n}\]So we need to solve for n, yes?

zepdrix (zepdrix):

Or equal to* woops.\[\Large 6\quad\ge\quad -\frac{4}{5n}\]